#8. Venkat is age 32, single, and reported AGI of $67,000 inntax last year 2018. He is an active participant in this enployeer's pension plan. What is the maximun deductible IRA contribution ge can make in 2018?

#9. Max retired in 2018 at age 62. During the year he received distribution of $9,000 from his IRA. He made nondeductible contributions of $20,000 to the IRA in prior years and has never received a nontaxable distribution. As of December 31, 2018, the value of his IRA was $150,000. Calculate the tax portion
Answer #1

8. The maximum deductible IRA contribution for the year 2018 = 5500*(67000-63000)/(73000-63000) = $2,200

9. The tax portion = 9000-9000*(20000/(150000+9000)) = $7,868